Badour Inc. is a job-order manufacturer. The company uses a predetermined overhead rate based on direct labor hours to apply overhead to individual jobs. For the current year, estimated direct labor hours were 126,000 and estimated factory overhead was $919,800. The following information was for September. Job X was completed during September, while Job Y was started but not finished. September 1, inventories: Materials $ 10,300 Work-in-process (All Job X) 38,600 Finished goods 81,600 Materials purchases $ 137,000 Direct materials requisitioned: Job X $ 55,700 Job Y 41,200 Direct labor hours: Job X 6,200 Job Y 5,700 Labor costs incurred: Direct labor ($7.20 per hour) $ 85,680 Indirect labor 17,400 Factory supervisory salaries 8,400 Rental costs: Factory $ 9,600 Administrative offices 3,400 Total equipment depreciation costs: Factory $ 10,300 Administrative offices 3,100 Indirect materials used $ 15,600 The total cost of Job X is: Multiple Choice $184,200. $141,300. $168,700. $135,600. $170,680.
Answer #1

Calculation of Total cost of X:

Work in process inventory = 38600

Direct materials of X= 55700

Direct labor (direct labor hours * rate per hour)= 6200*7.2= 44640

Overhead (overhead rate × labor hours)

Overhead rate= estimated factory overhead /estimated labor hours= 919800/126000= 7.3

Overhead = 7.3* 6200= 45260

Total cost = work in process inventory + direct materials + overhead =38600+55700+44640+45260

Total cost of X= 184200